OpenStudy (anonymous):

PLEASE HELP! Total cost of job = $147.51 Duration of job = 5 hours Price of parts = $32.50 Overhead rate = 55% Hourly labor rate = $______ per hour

OpenStudy (anonymous):

Do you need the hourly labor rate?

OpenStudy (anonymous):

I'm assuming by a0 you mean 10 or 20?

OpenStudy (anonymous):

NO we are trying to find the hourly rate and no i ment to put a blank!

OpenStudy (anonymous):

@adrianajones we are trying to find he hourly rate!

OpenStudy (anonymous):

$52.23 per hour

OpenStudy (phi):

the labor rate will not include the cost of parts. However, I do not know if the overhead rate is included or not. Do you know ? (see your book)

OpenStudy (anonymous):

Generally speaking I don't believe that overhead would be included in your labor costs, but like Phi said check your book. It will alter the outcome. So find the actual overhead price not including parts (Subtract parts from total cost) (total cost - parts = x) then x*0.55 (your 55% overhead) this will give you the overhead cost that we can remove from the total cost. Take your actual overhead price, subtract it from the total cost, then subtract materials. Divide that number by the number of hours worked.

OpenStudy (anonymous):

Its not 52.23 i already tried that and its a quiz... :( @Mike546 @phi @adrianajones

OpenStudy (anonymous):

I never said it was 52.23 if you follow my steps you get $10 and some change (deleted the answer already)

OpenStudy (anonymous):

Some one else said that and i tried that it was 10.23 or something!!

OpenStudy (anonymous):

Why is overhead rate not in dollars per time?

OpenStudy (anonymous):

If it's a quiz do you have certain options?

OpenStudy (anonymous):

Depends on if the question is from accounting or not. Then again I could be remembering something completely incorrectly....

OpenStudy (anonymous):

no its fill in the blank!!! THIS IS WHAT IT LOOKS LIKE 10.23 is not the anwser!

OpenStudy (anonymous):

*10.35

OpenStudy (anonymous):

@surjithayer Why are you multiplying the overhead rate by the cost of parts?

OpenStudy (anonymous):

let x be the hourly rate 5x+32.50=147.51 5x=147.51-32.50=115.01 x=115.01/5=23.002 overhead rate=23.002*.55=12.6511 hence hourly rate=23.002-12.6511=10.3509=10.35

OpenStudy (anonymous):

@surjithayer @Krisylynn22 attempted 10.35 and said it was not the answer

OpenStudy (anonymous):

it is not clear on which term overhead rate is applicable.

OpenStudy (anonymous):

it can be (5x+32.50)*155/100=147.51 5x+32.50=147.51*100/155=95.1677 5x=95.1677-32.50=62.6667 x=12.53

OpenStudy (anonymous):

Can you give us a formula from your textbook?